OpenStudy (anonymous):

c/1998 + d/1998 = 1/2 Find the values of c and d.

OpenStudy (anonymous):

infinite number of values for c and d

OpenStudy (mathteacher1729):

This is one equation in two unknowns. There are infinitely many values for c and d that make the equation true.

OpenStudy (anonymous):

\[c+d=999\]

OpenStudy (anonymous):

So whatever values you get, would they be marked correct? :O

OpenStudy (experimentx):

it's a line ... any point on this line would be your solution, and a line has infinite points

OpenStudy (mathteacher1729):

@emcrazy14 -- not just any values. Any values of c and d such that (c+d) / 1998 = 1/2

OpenStudy (anonymous):

Okay. I got c=998.001 and d=0.999. Are these correct?

OpenStudy (mathteacher1729):

Test those values by putting them into your original equation. :)

OpenStudy (experimentx):

try to get integer values, for best sake

OpenStudy (experimentx):

though that is also correct

OpenStudy (anonymous):

But how do I get the integer values? @experimentX

OpenStudy (experimentx):

put x=1, get y=998